### v2

# Higher curvature corrections to the black hole Wheeler-DeWitt equation and the annihilation to nothing scenario

Abstract
We revisit Yeom’s annihilation-to-nothing scenario using a modified Wheeler-DeWitt (WDW) equation, incorporating higher curvature corrections. By taking these corrections into account, we show that singularity resolution does not occur within low-energy effective field theory (EFT). Since general relativity (GR) is itself only a low-energy EFT of an underlying ultraviolet (UV) theory, it is unlikely that true singularity resolution can emerge within its domain of validity. Our analysis does not contradict Yeom’s conjecture but clarifies that any true resolution of the black hole singularity necessarily requires the inclusion of UV degrees of freedom beyond the scope of GR.
1 Introduction
The resolution of black hole singularities remains one of the central open problems in theoretical physics. General relativity (GR), despite its remarkable success as a classical theory of gravitation, inevitably predicts the formation of curvature singularities inside black holes [1, 2, 3]. Since such singularities signal the breakdown of the classical description, it is widely believed that a quantum theory of gravity is required to address this issue.
One of the well-studied approaches to the singularity problem is quantum cosmology based on the Wheeler-DeWitt (WDW) equation [4, 5, 6, 7]. Inside the black hole horizon, where the roles of the time and radial coordinates are interchanged, the geometry can be described by a minisuperspace model analogous to a homogeneous cosmology. The WDW equation is particularly suitable for this setting, as it naturally accommodates the time–radial coordinate interchange and allows a canonical analysis of quantum gravitational effects in the black hole interior [8, 9, 10, 11, 12, 13, 14]. Within this framework, Yeom and collaborators proposed the so-called Annihilation-to-nothing scenario [8, 9, 10, 11], in which two classical branches of spacetime—one evolving from the horizon and the other from the singularity—annihilate at a hypersurface inside the horizon. This interpretation provides a possible mechanism for the resolution of black hole singularities within the WDW formalism. However, whether such a mechanism survives once more fundamental quantum effects are included remains an open question.
In this broader context, nonperturbative approaches, such as loop quantum gravity [15, 16, 17, 18, 19, 20] or string theoretic constructions [21, 22], suggest that singularity resolution can occur due to fundamentally quantum effects. However, it is unclear whether such effects can be captured within a low energy effective field theory (EFT) framework. From the EFT perspective [23, 24, 25, 26], GR is viewed as the low energy limit of a more fundamental ultraviolet (UV) theory. Consequently, it is unnatural to expect that singularity resolution can be achieved solely within the canonical quantization of the GR framework, as its description inevitably breaks down near the Planck scale, where higher curvature operators and new physical degrees of freedom become significant.
In this paper, we revisit Yeom’s Annihilation-to-nothing interpretation by introducing higher-curvature corrections within the EFT framework, treating GR as an effective theory valid below a UV cutoff. By analyzing the modified WDW equation arising from this extended action, we re-evaluate Yeom’s wave packet solutions and their behavior in the presence of higher-curvature terms [27]. Our analysis clarifies that any genuine resolution of singularities cannot be achieved within the low-energy EFT framework and would require UV-sensitive operators beyond the GR regime.
The novelty of this paper is that we perform the first systematic EFT extension of the Wheeler–DeWitt equation for black hole interiors and show that Yeom’s annihilation-to-nothing feature does not survive once higher-curvature corrections are included. This reveals that the apparent singularity resolution in the classical WDW analysis cannot arise within low-energy effective theory.
The organization of this paper is as follows. In Sec. 2, we briefly review the WDW formulation inside black hole horizons and Yeom’s Annihilation-to-nothing interpretation. In Sec. 3, we discuss Yeom’s proposal by introducing higher-derivative corrections at the classical level within the EFT framework. In Sec. 4, we analyze the wave functions by incorporating the higher-derivative terms as quantum corrections in EFT and discuss their implications for Yeom’s interpretation. Finally, Sec. 5 presents our conclusion and outlook for future research.
Unless specified otherwise, we use the natural unit system, where the speed of light $c=1$ and the Dirac constant $\hbar=1$.
2 Review of Wheeler-DeWitt wave function inside black holes and Yeom’s interpretation
We review the derivation of the minisuperspace Wheeler DeWitt (WDW) equation inside the horizon of spherically symmetric black holes in arbitrary dimensions. According to Birkhoff’s theorem [28], any spherically symmetric vacuum solution of the Einstein equation must be static and is uniquely given by the Schwarzschild–Tangherlini metric in $D$ dimensions. Therefore, the interior region of a Schwarzschild black hole provides a representative setup for studying the minisuperspace dynamics. We further review Yeom’s interpretation of the WDW wave function for black hole interior in this section.
2.1 Wheeler-DeWitt equation inside a spherically symmetric black hole
We start from the Einstein-Hilbert action in D-dimensional spacetime,
$S=\frac{1}{16\pi G}\int_{M}d^{D}x\sqrt{-g}R,$ (2.1)
where $g$ is the determinant of the metric $g_{ab}$, $R$ is the D-dimensional Ricci scalar.
Inside the horizon, the Kantowski–Sachs–type metric is adopted[29],
$ds^{2}=-e^{2n(t)}dt^{2}+e^{2\alpha(t)}dr^{2}+r_{s}^{2}e^{2\beta(t)}d\Omega_{D-2}^{2},$ (2.2)
where $d\Omega_{D-2}^{2}$ is the line element of the (D2)(D-2)-sphere with unit radius and $r_{s}$ a constant of dimension length. This interior metric is related to the Schwarzschild–Tangherlini metric by
$e^{2\beta}=\frac{t^{2}}{r_{s}^{2}},\qquad e^{2\alpha}=\frac{r_{s}^{D-3}}{t^{D-3}}-1,\qquad e^{2n}=\!\left(\frac{r_{s}^{D-3}}{t^{D-3}}-1\right)^{-1},$ (2.3)
where time and radial coordinates are exchanged relative to the exterior region.
Choosing the lapse function as
$n(t)=\alpha(t)+(D-2)\beta(t),$ (2.4)
which can be achieved by a gauge transformation, and substituting the metric into the action, the minisuperspace reduction yields
$S\propto\!\int dt\!\left[-2\dot{\alpha}\dot{\beta}-(D-3)\dot{\beta}^{2}+(D-3)r_{s}^{2(D-3)}e^{2\alpha+2(D-3)\beta}\right].$ (2.5)
The action can further be rewritten in terms of the variables $X=\alpha$ and $Y=\alpha+(D-3)\beta$ as
$S\propto\!\int dt\!\left[\frac{1}{D-3}(\dot{X}^{2}-\dot{Y}^{2})+(D-3)r_{s}^{2(D-3)}e^{2Y}\right].$ (2.6)
Therefore, we obtain the canonical Hamiltonian
$H=(D-3)\left[\frac{1}{4}(\Pi_{X}^{2}-\Pi_{Y}^{2})-r_{s}^{2(D-3)}e^{2Y}\right],$ (2.7)
where $\Pi_{X}$ and $\Pi_{Y}$ are the canonical conjugate momenta to $X$ and $Y$, respectively. Through canonical quantization, the canonical momenta are replaced by differential operators acting on the wave function,
$\Pi_{X}\rightarrow\hat{\Pi}_{X}=-i\frac{\partial}{\partial X},\qquad\Pi_{Y}\rightarrow\hat{\Pi}_{Y}=-i\frac{\partial}{\partial Y},$ (2.8)
resulting in the Wheeler–DeWitt (WDW) equation in the minisuperspace [4, 5, 6, 7]. The WDW equation arises from the Hamiltonian constraint $\hat{\mathcal{H}}\Psi=0$, which requires that the wave function of the universe satisfy this constraint in quantum gravity. Thus, the minisuperspace WDW equation can be written as
$\left(\frac{\partial^{2}}{\partial X^{2}}-\frac{\partial^{2}}{\partial Y^{2}}+4r_{s}^{2(D-3)}e^{2Y}\right)\Psi(X,Y)=0.$ (2.9)
This form is dimension-independent up to the power of $r_{s}$.
In this coordinate system, the region $X,Y\rightarrow$-$\infty$ represents the event horizon, whereas the limit $X\rightarrow\infty$ and $Y\rightarrow$-$\infty$ corresponds to the spacetime singularity.
2.2 Solutions and Yeom’s interpretation
The fundamental solution of (2.9) is
$\psi_{k}(X,Y)=e^{-ikX}K_{ik}(2r_{s}e^{Y}),$ (2.10)
and general solutions are given by
$\Psi(X,Y)=\int_{-\infty}^{\infty}f(k)\,\psi_{k}(X,Y)\,dk,$ (2.11)
where $f(k)$ is an amplitude function.
Using the asymptotic form of the modified Bessel function, one can construct Gaussian or analytical wave packets. For example, choosing $f(k)=ik$ yields
$\Psi_{1}(X,Y)=2\pi r_{s}^{D-3}e^{Y}\sinh X\,e^{-2r_{s}^{D-3}e^{Y}\cosh X}.$ (2.12)
The probability density $\rho=|\Psi_{1}|^{2}$ shows a peak following the classical trajectory
$e^{Y}\cosh X=\mathrm{const.},$ (2.13)
and it vanishes near $X=0$, where neither the horizon nor the singularity exists. This behavior is interpreted as the “Annihilation-to-nothing” process proposed by Yeom et al. [8, 9, 10, 11], representing the annihilation of two classical spacetime branches with opposite time arrows inside the horizon. The behavior of the probability density on XYX-Y plane is shown in Fig. 1.
Fig. 1: Probability density of the wave function for $r_{s}=1$. Adapted from Ref. [8].
To summarize this section, the WDW wave packet inside spherical black holes exhibits behavior consistent with an annihilation-to-nothing process, thereby supporting Yeom’s interpretation. Related analyses have also been carried out for topological black holes using the scalar-type WDW equation and for black holes described by the Dirac-type WDW equation [30, 31, 32]. For topological black holes, the results remain compatible with Yeom’s interpretation, whereas the Dirac-type formulation yields a novel variant of the annihilation picture.
General relativity is regarded as a low-energy effective theory and is unlikely to capture the nonperturbative effects relevant to singularity resolution. Thus, the apparent resolution observed in the above analyses likely arises from the fact that higher-derivative corrections— which generically appear in the effective field theory of gravity—have not been incorporated. In the following sections, we examine this issue in detail: in Sec. 3 we introduce classical higher-derivative corrections, while in Sec. 4 we analyze quantum effects arising from curvature-squared terms that vanish at the classical level. These analyses provide new insights into Yeom’s interpretation.
3 Wheeler-DeWitt wave function with classical higher curvature corrections
In this section, we consider classical corrections to the WDW wave function in the black-hole interior within the framework of effective field theory. Since we focus on vacuum solutions, the Ricci tensor vanishes identically, and the action can therefore be constructed solely from the Riemann tensor. Modulo tensor identities [33], the only independent fourth-derivative invariant that does not involve the Ricci tensor, is $R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$; this is field-redefinition equivalent to the Gauss–Bonnet combination $R^{2}-4R_{\mu\nu}R^{\mu\nu}+R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$, which in four dimensions is topological and therefore does not affect the bulk equations of motion. At the next order the independent Riemann-constructed operators are $\nabla_{\alpha}R_{\mu\nu\rho\sigma}\nabla^{\alpha}R^{\mu\nu\rho\sigma}$ and $R^{\rho\sigma}_{\ \ \mu\nu}R^{\alpha\beta}_{\ \ \rho\sigma}R^{\mu\nu}_{\ \ \alpha\beta}$; the former can be rearranged, up to a total derivative, into a Riemann-cubed term plus Ricci-dependent operators and so can be dropped without loss of generality. Likewise, there are two independent components of the fourth power of Riemann tensors. Thus, we consider the following action in four dimensional spacetime,
$S=\frac{1}{16\pi G}\int d^{4}x\sqrt{-g}(R+\gamma R^{\rho\sigma}_{\ \ \mu\nu}R^{\alpha\beta}_{\ \ \rho\sigma}R^{\mu\nu}_{\ \ \alpha\beta}+\eta\mathcal{C}^{2}+\lambda\tilde{\mathcal{C}}^{2}),$ (3.1)
where $\mathcal{C}$ and $\tilde{\mathcal{C}}$ are defined by
$\mathcal{C}$ $=R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma},$ (3.2)
$\tilde{\mathcal{C}}$ $=\tilde{R}_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma},$ (3.3)
with the dual tensor $\tilde{R}_{\mu\nu\rho\sigma}=\epsilon_{\mu\nu}^{\ \ \alpha\beta}R_{\alpha\beta\rho\sigma}$.
As discussed in the previous section, the metric inside a black hole can be written in the Kantowski–Sachs form,
$ds^{2}=g_{\mu\nu}dx^{\mu}dx^{\nu}=-e^{2\alpha(t)+4\beta(t)}dt^{2}+e^{2\alpha(t)}dr^{2}+r_{s}^{2}e^{2\beta(t)}d\Omega_{2}^{2}.$ (3.4)
For the Schwarzschild spacetime, the functions take the specific form
$e^{2\alpha(t)}=\frac{r_{s}}{\tau(t)}-1,\ \ \ e^{2\beta(t)}=\frac{\tau^{2}(t)}{r_{s}^{2}}.$ (3.5)
Note that $\tau$ denotes the radial coordinate of the familiar Schwarzschild metric, and in the present case it is taken to be a function of the spacetime coordinate $t$. Then, the higher derivative corrections are
$R^{\rho\sigma}_{\ \ \mu\nu}R^{\alpha\beta}_{\ \ \rho\sigma}R^{\mu\nu}_{\ \ \alpha\beta}=\frac{12r_{s}^{3}}{\tau^{9}(t)},\ \ \ \mathcal{C}^{2}=\frac{144r_{s}^{4}}{\tau^{12}(t)},\ \ \ \tilde{\mathcal{C}}^{2}=0.$ (3.6)
Using the variables $X$ and $Y$ introduced in the previous section, these are replaced with the following operators in the quantum Hamiltonian constraint, $\hat{\mathcal{H}}\Psi=0$.
$R^{\rho\sigma}_{\ \ \mu\nu}R^{\alpha\beta}_{\ \ \rho\sigma}R^{\mu\nu}_{\ \ \alpha\beta}\rightarrow\frac{12}{r_{s}^{6}}e^{9X-9Y},\ \ \ \mathcal{C}^{2}\rightarrow\frac{144}{r_{s}^{8}}e^{12X-12Y},\ \ \ \tilde{\mathcal{C}}^{2}\rightarrow 0.$ (3.7)
The WDW equation then takes the following form.
$\hat{\mathcal{H}}\Psi=-\left(\frac{1}{4}\frac{\partial^{2}}{\partial X^{2}}-\frac{1}{4}\frac{\partial^{2}}{\partial Y^{2}}+r_{s}^{2}e^{2\hat{Y}}-\frac{12\gamma}{r_{s}^{4}}e^{7X-5Y}-\frac{144\eta}{r_{s}^{6}}e^{10X-8Y}\right)\Psi(X,Y)$
$\qquad=0.$ (3.8)
Expanding the WDW equation perturbatively in terms of the higher curvature couplings $\gamma$ and $\eta$, we separate the zeroth- and first-order contributions as follows:
$\left(\frac{1}{4}\frac{\partial^{2}}{\partial X^{2}}-\frac{1}{4}\frac{\partial^{2}}{\partial Y^{2}}+r_{s}^{2}e^{2\hat{Y}}\right)\Psi^{(0)}(X,Y)=0,$ (3.9)
$\left(\frac{1}{4}\frac{\partial^{2}}{\partial X^{2}}-\frac{1}{4}\frac{\partial^{2}}{\partial Y^{2}}+r_{s}^{2}e^{2\hat{Y}}\right)\Psi^{(1)}(X,Y)=\left(\frac{12\gamma}{r_{s}^{4}}e^{7X-5Y}+\frac{144\eta}{r_{s}^{6}}e^{10X-8Y}\right)\Psi^{(0)}(X,Y).$ (3.10)
To facilitate the analysis, we introduce a change of variables $(X,Y)\rightarrow(Z,W)$ such that the WDW operator, including both the kinetic and potential terms, is brought to a standard form up to an overall conformal factor. In terms of these new variables, the operator takes the form
$\left(\frac{1}{4}\frac{\partial^{2}}{\partial X^{2}}-\frac{1}{4}\frac{\partial^{2}}{\partial Y^{2}}+r_{s}^{2}e^{2\hat{Y}}\right)=(W^{2}-Z^{2})\left(\frac{1}{4}\frac{\partial^{2}}{\partial Z^{2}}-\frac{1}{4}\frac{\partial^{2}}{\partial W^{2}}-1\right),$ (3.11)
where the new coordinates are defined by
$Z=r_{s}e^{Y}\cosh X,\ \ \ \ W=r_{s}e^{Y}\sinh X,$ (3.12)
which can be inverted as
$e^{Y}=\frac{\sqrt{Z^{2}-W^{2}}}{r_{s}},\ \ \ e^{X}=\frac{Z+W}{\sqrt{Z^{2}-W^{2}}}.$ (3.13)
This transformation maps the minisuperspace into the wedge defined by $Z>|W|$.
Substituting these relations into Eq. (3.10), the first-order wave function $\Psi^{(1)}$ can be expressed in terms of the Green’s function $G(Z,W;Z^{\prime},W^{\prime})$ of the two-dimensional Klein–Gordon operator as
$\Psi^{(1)}(Z,W)=-\int\!\!\int_{Z^{\prime}>|W^{\prime}|,Z^{\prime}\geq 0}dZ^{\prime}dW^{\prime}G(Z,W;Z^{\prime},W^{\prime})\frac{1}{Z^{\prime 2}-W^{\prime 2}}$
$\qquad\qquad\qquad\qquad\qquad\qquad\times\left(\frac{12\gamma}{r_{s}^{4}}e^{7X^{\prime}-5Y^{\prime}}+\frac{144\eta}{r_{s}^{6}}e^{10X^{\prime}-8Y^{\prime}}\right)\Psi^{(0)}(Z^{\prime},W^{\prime}),$ (3.14)
where the Green’s function is given by
$G(Z,W;Z^{\prime},W^{\prime})$ $=\int\frac{d^{2}k}{\pi^{2}}\frac{e^{-i\left(k_{2}(W-W^{\prime})-k_{1}(Z-Z^{\prime})\right)}}{k_{2}^{2}-k_{1}^{2}-4},$ (3.15)
$=-\theta(s)H_{0}^{(2)}(2\sqrt{s})-\frac{2i}{\pi}\theta(-s)K_{0}(2\sqrt{-s}),$ (3.16)
where $H_{0}^{(2)}(x)$ is the Hankel function of second kind, $K_{0}(x)$ is the modified Bessel function and $s=(W-W^{\prime})^{2}-(Z-Z^{\prime})^{2}$ for the interval [34, 35]. The Green function satisfies the equation
$\left(\frac{1}{4}\frac{\partial^{2}}{\partial Z^{2}}-\frac{1}{4}\frac{\partial^{2}}{\partial W^{2}}-1\right)G(Z,W;Z^{\prime},W^{\prime})=\delta(Z-Z^{\prime})\delta(W-W^{\prime}).$ (3.17)
After inserting the explicit form of $\Psi^{(0)}$ and changing variables, this integral becomes
$\Psi^{(1)}(Z,W)$ $=-24\pi r_{s}\int\!\!\int_{Z^{\prime}>|W^{\prime}|}\frac{dZ^{\prime}dW^{\prime}}{Z^{\prime 2}-W^{\prime 2}}G(Z,W;Z^{\prime},W^{\prime})\left(\frac{\gamma}{r_{s}^{4}}e^{7X^{\prime}-5Y^{\prime}}+\frac{12\eta}{r_{s}^{6}}e^{10X^{\prime}-8Y^{\prime}}\right)$
$\qquad\qquad\qquad\qquad\times e^{Y^{\prime}}\sinh X^{\prime}e^{-2r_{s}e^{Y\prime}\cosh X^{\prime}}$
$=-24\pi r_{s}\int\!\!\int_{Z^{\prime}>|W^{\prime}|}dZ^{\prime}dW^{\prime}G(Z,W;Z^{\prime},W^{\prime})\frac{W^{\prime}(Z^{\prime}+W^{\prime})^{7}}{(Z^{\prime 2}-W^{\prime 2})^{7}}\left(\gamma+12\frac{\eta r_{s}(Z^{\prime}+W^{\prime})^{3}}{(Z^{\prime 2}-W^{\prime 2})^{3}}\right)$
$\qquad\qquad\qquad\qquad\times e^{-2Z^{\prime}}.$ (3.18)
The solution to the Wheeler–DeWitt (WDW) equation is highly sensitive to the choice of boundary conditions. In this work, we adopt the Feynman boundary condition as the physically appropriate prescription. The perturbative construction of the WDW solution satisfying this condition was discussed in Ref. [36].
We aim to investigate whether the wave function continues to vanish at $X=0$ once higher-curvature corrections are incorporated within the EFT framework. Since singularity resolution is generally expected to originate from genuinely nonperturbative quantum gravitational effects, such behavior is unlikely to arise in a low-energy effective theory. In what follows, we analyze the behavior of the wave function in the vicinity of $X=0$.
For clarity of presentation, we omit the detailed algebra and present the full derivation in Appendix A. The perturbative WDW wave function in the region $X\approx 0$ and Y<<1Y<<-1 is given by
$\Psi^{(1)}\big|_{X=0,\,Y\ll-1}=-\frac{16777216\sqrt{\pi}\,r_{s}\gamma}{e^{8}}\Big(\frac{\ln\epsilon}{\epsilon^{12}}+\frac{1}{12}\frac{1}{\epsilon^{12}}+\frac{\gamma_{E}}{\epsilon^{12}}-\frac{\ln a}{a^{12}}-\frac{1}{12}\frac{1}{a^{12}}-\frac{\gamma_{E}}{a^{12}}\Big)$
$\qquad\qquad\qquad\qquad+\frac{414998793616\sqrt{22\pi}\,r_{s}^{2}\eta}{e^{11}}\Big(\frac{\ln\epsilon}{\epsilon^{18}}+\frac{1}{18}\frac{1}{\epsilon^{18}}+\frac{\gamma_{E}}{\epsilon^{18}}-\frac{\ln a}{a^{18}}-\frac{1}{18}\frac{1}{a^{18}}-\frac{\gamma_{E}}{a^{18}}\Big).$
Here, $\epsilon$ denotes the lower cutoff of the $\rho$-integration, while $a$ represents the upper cutoff chosen to delimit the region that gives the dominant contribution to the integral. With these cutoffs, the leading small-$\epsilon$ behavior is dominated by $\epsilon^{-12}\ln\epsilon$ and $\epsilon^{-18}\ln\epsilon$ terms coming from the two contributions respectively.
Then, the wave function implies
$\Psi^{(1)}\big|_{X=0,\,Y\ll-1}\neq 0,$ (3.20)
and although $\Psi^{(1)}$ exhibits divergent behavior as $\epsilon\to 0$ with the above leading scalings, these divergent contributions are expected to be removed by an appropriate regularization procedure, yielding a finite result.
Substituting these results into the expression for the first-order correction, we obtain the following non-vanishing contribution to the probability density of the wave function $\rho(X,Y)$ in the vicinity of $X\approx 0$ and $Y\ll-1$:
$\rho(0,Y)=|\Psi(0,Y)|^{2}_{Y\ll-1}$ $=|\Psi^{(0)}(0,Y)+\Psi^{(1)}(0,Y)|^{2}_{Y\ll-1}$
$=|\Psi^{(1)}(0,Y)|^{2}_{Y\ll-1}$
$\neq 0.$ (3.21)
By considering the first-order perturbative correction, we find that the wave function of the universe does not vanish around $X=0$ and $Y\ll-1$. For computational simplicity, we have evaluated the expression in the regime $Y\ll-1$ with appropriate regularization. Although the precise value depends on the chosen regularization, it is expected that the wave function remains nonzero for generic values of $Y$. This indicates that within the low-energy effective theory, the classical singularity is not resolved, and we therefore claim that its resolution requires a UV-complete theory of quantum gravity.
4 Wheeler-DeWitt wave function for quantum higher curvature corrections
In the previous section we examined classical higher-curvature corrections to the WDW wave function in the black-hole interior within the framework of effective field theory. This analysis indicates that, within the regime of validity of the low-energy effective theory, the classical singularity remains unresolved, suggesting that its resolution requires a UV-complete theory of quantum gravity. In this section we consider four-dimensional quadratic gravity as an effective field theory and analyze how quantum corrections modify the WDW wave function in this framework. The action of quadratic gravity is given by
$S=\int d^{4}x\sqrt{-g}(R+\gamma R^{2}+\eta R_{\mu\nu}R^{\mu\nu}+\lambda E_{4}),$ (4.1)
where $E_{4}$ is the topological Euler term,
$E_{4}=R^{2}-4R_{\mu\nu}R^{\mu\nu}+R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}.$ (4.2)
In four-dimensional spacetime, this topological term becomes a total derivative and thus does not contribute to the local equations of motion. The physical dynamics are therefore governed by the Ricci scalar and Ricci tensor terms.
Inside the event horizon, we consider a Kantowski–Sachs–type metric, following the same procedure as in the previous section,
$ds^{2}=g_{\mu\nu}dx^{\mu}dx^{\nu}=-e^{2\alpha(t)+4\beta(t)}dt^{2}+e^{2\alpha(t)}dr^{2}+r_{s}^{2}e^{2\beta(t)}d\Omega_{2}^{2}.$ (4.3)
By substituting this metric into the quadratic gravity action, we obtain
$S=\int d^{4}x\Bigl[-2\dot{\alpha}\dot{\beta}-\dot{\beta}^{2}+r_{s}^{2}e^{2\alpha+2\beta}+\frac{\gamma}{r_{s}^{4}}(-2\dot{\alpha}\dot{\beta}-\dot{\beta}^{2}+r_{s}^{2}e^{2\alpha+2\beta}+\ddot{\alpha}+2\ddot{\beta})^{2}e^{-2(\alpha+2\beta)}$
$\qquad\qquad\ \ +\frac{\eta}{r_{s}^{4}}\Bigl(\ddot{\alpha}+(4\dot{\alpha}\dot{\beta}+2\dot{\beta}^{2}-\ddot{\alpha}-2\ddot{\beta})^{2}+2(\ddot{\beta}+r_{s}^{2}e^{2Y})^{2}\Bigr)e^{-2(\alpha+2\beta)}\Bigr].$ (4.4)
To simplify the expression, it is convenient to introduce the field redefinitions $X(t)=\alpha(t)$ and $Y(t)=\alpha(t)+\beta(t)$, as in the previous section. In terms of these new variables, the action takes the form
$S=\int d^{4}x\Bigl[\dot{X}^{2}-\dot{Y}^{2}+r_{s}^{2}e^{2Y}+\frac{\gamma}{r_{s}^{4}}e^{2X-4Y}(\dot{X}^{2}-\dot{Y}^{2}+r_{s}^{2}e^{2Y}-\ddot{X}+2\ddot{Y})^{2}$
$\qquad\qquad\ \ +\frac{\eta}{r_{s}^{4}}e^{2X-4Y}\Bigl(\ddot{X}^{2}+(-2\dot{X}^{2}+2\dot{Y}^{2}+\ddot{X}-2\ddot{Y})^{2}+2(-\ddot{X}+\ddot{Y}+r_{s}^{2}e^{2Y})^{2}\Bigr)\Bigr].$ (4.5)
In this work, we basically employ the perturbative analysis with small parameter $\gamma,\eta=\mathcal{O}(\epsilon)$. In this Lagrangean, the motion fo equations on the zeroth order are
$\qquad\ \ \ \ \ \ddot{X}=0,$ (4.6)
$\ddot{Y}+r_{s}^{2}e^{2Y}=0,$ (4.7)
which reproduce the Einstein gravity limit without higher-derivative effects.
In the framework of effective field theory, no new physical degrees of freedom, which are absent in Einstein gravity, appear at this perturbative order. To remove spurious higher-derivative terms, we perform field redefinitions as
$X\rightarrow X-\frac{\gamma}{2r_{s}^{4}}e^{2X-4Y}\left(2\dot{X}^{2}-2\dot{Y}^{2}+2r_{s}^{2}e^{2Y}-\ddot{X}+3\ddot{Y}\right)$
$\ \ \ \ \ \ \ -\frac{2\eta}{r_{s}^{4}}e^{2X-4Y}(\dot{X}^{2}-\dot{Y}^{2}-\ddot{X}+\ddot{Y}),$ (4.8)
$Y\rightarrow Y-\frac{2\gamma}{r_{s}^{4}}e^{2X-4Y}(\dot{X}^{2}-\dot{Y}^{2}+\ddot{Y})-\frac{\eta}{r_{s}^{2}}(4\dot{X}^{2}-4\dot{Y}^{2}-2\ddot{X}+3\ddot{Y}-r_{s}^{2}e^{2Y}).$ (4.9)
These transformations eliminate all second-derivative contributions from the Lagrangian, yielding an equivalent, low-energy description in which ghost-like excitations are integrated out.
Thus, the resulting simplified action becomes
$S=\int d^{4}x\Bigl[\dot{X}^{2}-\dot{Y}^{2}+r_{s}^{2}e^{2Y}+\frac{\gamma}{r_{s}^{4}}e^{2X-4Y}(\dot{X}^{2}-\dot{Y}^{2}+r_{s}^{2}e^{2Y})^{2}-\frac{4\gamma}{r_{s}^{2}}e^{2X-2Y}(\dot{X}^{2}-\dot{Y}^{2})$
$\qquad\qquad\ \ +\frac{4\eta}{r_{s}^{4}}e^{2X-4Y}\left(\dot{X}^{2}-\dot{Y}^{2}\right)\left(\dot{X}^{2}-\dot{Y}^{2}-2r_{s}^{2}e^{2Y}\right)+4\eta e^{2X}\Bigr]$
$\ \ \ =\int d^{4}xL(\dot{X},X,\dot{Y},Y).$ (4.10)
4.1 Hamiltonian constraint and Wheeler-DeWitt wave function
Let us perform the Hamiltonian analysis by considering this Lagrangian, which is physically equivalent to the original one. From this action, we can compute the conjugate momenta
$\Pi_{X}=2\dot{X}+\frac{2\gamma}{r_{s}^{4}}e^{2X-4Y}\Pi_{X}\left(\frac{1}{4}\Pi_{X}^{2}-\frac{1}{4}\Pi_{Y}^{2}-r_{s}^{2}e^{2Y}\right)$
$\qquad\ \ \ +\frac{8\eta}{r_{s}^{4}}e^{2X-4Y}\Pi_{X}\left(\frac{1}{4}\Pi_{X}^{2}-\frac{1}{4}\Pi_{Y}^{2}-r_{s}^{2}e^{2Y}\right)+\mathcal{O}(\epsilon^{2}),$ (4.11)
$\Pi_{Y}=-2\dot{Y}+\frac{2\gamma}{r_{s}^{4}}e^{2X-4Y}\Pi_{Y}\left(\frac{1}{4}\Pi_{X}^{2}-\frac{1}{4}\Pi_{Y}^{2}-r_{s}^{2}e^{2Y}\right)$
$\qquad\ \ \ +\frac{8\eta}{r_{s}^{4}}e^{2X-4Y}\Pi_{Y}\left(\frac{1}{4}\Pi_{X}^{2}-\frac{1}{4}\Pi_{Y}^{2}-r_{s}^{2}e^{2Y}\right)+\mathcal{O}(\epsilon^{2}).$ (4.12)
Then, the Hamiltonian density is
$H=\int dx^{3}\Bigl[\Pi_{X}\dot{X}+\Pi_{Y}\dot{Y}-L(X,\dot{X},Y,\dot{Y})\Bigr]$
$\ \ \ =\int dx^{3}\Bigl[\frac{1}{4}\Pi_{X}^{2}-\frac{1}{4}\Pi_{Y}^{2}-r_{s}^{2}e^{2Y}-\frac{1}{r_{s}^{4}}(\gamma+4\eta)\left(\frac{1}{4}\Pi_{X}^{2}-\frac{1}{4}\Pi_{Y}^{2}-r_{s}^{2}e^{2Y}\right)^{2}e^{2X-4Y}\Bigr].$ (4.13)
The Hamiltonian can be written in terms of the Hamiltonian and momentum constraints, from which the WDW equation is obtained. In the canonical quantization scheme, this equation arises from promoting the Hamiltonian constraint to a quantum operator. The canonical formulation of four-dimensional quadratic gravity, including the explicit forms of these constraints, has been discussed in detail in Refs. [37, 38, 39].
In the next step, using the Weyl ordering, we replace the classical observables with quantum operators. In this prescription, a classical observable $A(q,p)$ is promoted to an operator according to
$\hat{A}(\hat{q},\hat{\Pi})=e^{\frac{1}{2i}\frac{\partial}{\partial q}\frac{\partial}{\partial\Pi}}A_{0}(q,\Pi)\mid_{q=\hat{q},\Pi=\hat{\Pi}}$. (4.14)
By expandong the second term of Eq.(4.1), we find that:
$\ \ \ \left(\frac{1}{4}\Pi_{X}^{2}-\frac{1}{4}\Pi_{Y}^{2}-r_{s}^{2}e^{2Y}\right)^{2}e^{2X-4Y}$
$=\frac{1}{16}\Pi_{X}^{4}e^{2X-4Y}+\frac{1}{16}\Pi_{Y}^{2}e^{2X-4Y}+r_{s}^{4}e^{2X}-\frac{1}{8}\Pi_{X}^{2}\Pi_{Y}^{2}e^{2X-4Y}$
$\ \ \ -\frac{1}{2}r_{s}^{2}\Pi_{X}^{2}e^{2X-2Y}+\frac{1}{2}\Pi_{Y}^{2}r_{s}^{2}e^{2X-2Y}.$ (4.15)
Each term in the right-hand side is then quantized according to the Weyl ordering as follows.
$\Pi_{X}^{4}e^{2X-4Y}\rightarrow e^{2\hat{X}-4\hat{Y}}\hat{\Pi}_{X}^{4}+\frac{4}{i}e^{2\hat{X}-4\hat{Y}}\hat{\Pi}_{X}^{3}-6e^{2\hat{X}-4\hat{Y}}\hat{\Pi}_{X}^{2}$
$\qquad\qquad\ \ \ \ \ \ -\frac{4}{i}e^{2\hat{X}-4\hat{Y}}\hat{\Pi}_{X}+e^{2\hat{X}-4\hat{Y}}$, (4.16)
$\Pi_{Y}^{4}e^{2X-4Y}\rightarrow e^{2\hat{X}-4\hat{Y}}\hat{\Pi}_{Y}^{4}-\frac{8}{i}e^{2\hat{X}-4\hat{Y}}\hat{\Pi}_{Y}^{3}-24e^{2\hat{X}-4\hat{Y}}\hat{\Pi}_{Y}^{2}$
$\qquad\qquad\ \ \ \ \ \ +\frac{32}{i}e^{2\hat{X}-4\hat{Y}}\hat{\Pi}_{Y}+16e^{2\hat{X}-4\hat{Y}}$, (4.17)
$\Pi_{X}^{2}\Pi_{Y}^{2}e^{2X-4Y}\rightarrow e^{2\hat{X}-4\hat{Y}}\hat{\Pi}_{X}^{2}\hat{\Pi}_{Y}^{2}-\frac{4}{i}e^{2\hat{X}-4\hat{Y}}\hat{\Pi}_{X}^{2}\hat{\Pi}_{Y}-4e^{2\hat{X}-4\hat{Y}}\hat{\Pi}_{X}^{2}$
$\qquad\qquad\ \ \ \ \ \ +\frac{2}{i}e^{2\hat{X}-4\hat{Y}}\hat{\Pi}_{X}\hat{\Pi}_{Y}^{2}+8e^{2\hat{X}-4\hat{Y}}\hat{\Pi}_{X}\hat{\Pi}_{Y}-\frac{8}{i}e^{2\hat{X}-4\hat{Y}}\hat{\Pi}_{X}$
$\qquad\qquad\ \ \ \ \ \ -e^{2\hat{X}-4\hat{Y}}\hat{\Pi}_{Y}^{2}+\frac{4}{i}e^{2\hat{X}-4\hat{Y}}\hat{\Pi}_{Y}+4e^{2\hat{X}-4\hat{Y}}$ (4.18)
$\Pi_{X}^{2}e^{2X-2Y}\rightarrow e^{2\hat{X}-2\hat{Y}}\hat{\Pi}_{X}^{2}+\frac{2}{i}e^{2\hat{X}-2\hat{Y}}\hat{\Pi}_{X}-e^{2\hat{X}-2\hat{Y}}$, (4.19)
$\Pi_{Y}^{2}e^{2X-2Y}\rightarrow e^{2\hat{X}-2\hat{Y}}\hat{\Pi}_{Y}^{2}-\frac{2}{i}e^{2\hat{X}-2\hat{Y}}\hat{\Pi}_{Y}-e^{2\hat{X}-2\hat{Y}}$. (4.20)
Because the spacetime under consideration is static, the total Hamiltonian becomes proportional to the Hamiltonian constraint. Then, the quantum Hamiltonian constraint $\hat{\mathcal{H}}$ can be written in terms of quantum operators as
$\hat{\mathcal{H}}=\left(\frac{1}{4}\frac{\partial^{2}}{\partial X^{2}}-\frac{1}{4}\frac{\partial^{2}}{\partial Y^{2}}+r_{s}^{2}e^{2\hat{Y}}\right)+(\gamma+4\eta)e^{2\hat{X}}$
$\qquad\ +\frac{1}{4r_{s}^{2}}(\gamma+4\eta)e^{2\hat{X}-2\hat{Y}}\Bigl(\frac{\partial^{2}}{\partial X^{2}}+\frac{\partial}{\partial X}+1\Bigr)$
$\qquad\ +\frac{1}{16r_{s}^{4}}(\gamma+4\eta)e^{2\hat{X}-4\hat{Y}}\Bigl(\frac{\partial^{4}}{\partial X^{4}}+4\frac{\partial^{3}}{\partial X^{3}}+6\frac{\partial^{2}}{\partial X^{2}}+4\frac{\partial}{\partial X}+1\Bigr)$
$\qquad\ -\frac{1}{16r_{s}^{4}}(\gamma+4\eta)e^{2\hat{X}-4\hat{Y}}\Bigl(\frac{\partial^{4}}{\partial X^{2}\partial Y^{2}}-4\frac{\partial^{3}}{\partial X^{2}\partial Y}+\frac{\partial^{2}}{\partial X^{2}}+2\frac{\partial^{3}}{\partial X\partial Y^{2}}$
$\qquad\ -8\frac{\partial^{2}}{\partial X\partial Y}+8\frac{\partial}{\partial X}+\frac{\partial^{2}}{\partial Y^{2}}-4\frac{\partial}{\partial Y}+4\Bigr)$
$\qquad\ -\frac{1}{4r_{s}^{2}}(\gamma+4\eta)e^{2\hat{X}-2\hat{Y}}\Bigl(\frac{\partial^{2}}{\partial Y^{2}}-\frac{\partial}{\partial Y}+1\Bigr)$
$\qquad\ +\frac{1}{16r_{s}^{4}}(\gamma+4\eta)e^{2\hat{X}-4\hat{Y}}\Bigl(\frac{\partial^{4}}{\partial Y^{4}}-8\frac{\partial^{3}}{\partial Y^{3}}+24\frac{\partial^{2}}{\partial Y^{2}}-32\frac{\partial}{\partial Y}+16\Bigr),$ (4.21)
where $\hat{\Pi}_{X}=-i\frac{\partial}{\partial X}$ and $\hat{\Pi}_{Y}=-i\frac{\partial}{\partial Y}$.
Now, the Hamiltonian constraint derives the perturbative WDW equation for the wave function $\Psi(X,Y)$,
$\hat{\mathcal{H}}^{(0)}\Psi^{(1)}(X,Y)=-\hat{\mathcal{H}}^{(1)}\Psi^{(0)}(X,Y),$ (4.22)
where the zeroth- and first-order contributions to the Hamiltonian operator are denoted by $\hat{\mathcal{H}}^{(0)}$ and $\hat{\mathcal{H}}^{(1)}$, respectively. Here, $\hat{\mathcal{H}}^{(0)}$ represents the leading-order Hamiltonian, corresponding to Einstein gravity, while $\hat{\mathcal{H}}^{(1)}$ encodes the quadratic-curvature corrections arising from the higher-derivative terms.
[…]
$\hat{\mathcal{H}}^{(1)}=(\gamma+4\eta)e^{2\hat{Y}}+\frac{1}{4r_{s}^{2}}(\gamma+4\eta)e^{2\hat{X}-2\hat{Y}}\Bigl(\frac{\partial^{2}}{\partial X^{2}}+\frac{\partial}{\partial X}+1\Bigr)$
$\qquad\ \ \ \ +\frac{1}{16r_{s}^{4}}(\gamma+4\eta)e^{2\hat{X}-4\hat{Y}}\Bigl(\frac{\partial^{4}}{\partial X^{4}}+4\frac{\partial^{3}}{\partial X^{3}}+6\frac{\partial^{2}}{\partial X^{2}}+4\frac{\partial}{\partial X}+1\Bigr)$
$\qquad\ \ \ \ -\frac{1}{16r_{s}^{4}}(\gamma+4\eta)e^{2\hat{X}-4\hat{Y}}\Bigl(\frac{\partial^{4}}{\partial X^{2}\partial Y^{2}}-4\frac{\partial^{3}}{\partial X^{2}\partial Y}+\frac{\partial^{2}}{\partial X^{2}}+2\frac{\partial^{3}}{\partial X\partial Y^{2}}$
$\qquad\ \ \ \ -8\frac{\partial^{2}}{\partial X\partial Y}+8\frac{\partial}{\partial X}+\frac{\partial^{2}}{\partial Y^{2}}-4\frac{\partial}{\partial Y}+4\Bigr)$
$\qquad\ \ \ \ -\frac{1}{4r_{s}^{2}}(\gamma+4\eta)e^{2\hat{X}-2\hat{Y}}\Bigl(\frac{\partial^{2}}{\partial Y^{2}}-\frac{\partial}{\partial Y}+1\Bigr)$
$\qquad\ \ \ \ +\frac{1}{16r_{s}^{4}}(\gamma+4\eta)e^{2\hat{X}-4\hat{Y}}\Bigl(\frac{\partial^{4}}{\partial Y^{4}}-8\frac{\partial^{3}}{\partial Y^{3}}+24\frac{\partial^{2}}{\partial Y^{2}}-32\frac{\partial}{\partial Y}+16\Bigr).$ (4.24)
As in the previous section, the first-order correction to the wave function, $\Psi^{(1)}(Z,W)$, can be expressed formally in terms of the Green’s function (3.15) as follows:
$\Psi^{(1)}(Z,W)=-\int\!\!\int_{Z^{\prime}>|W^{\prime}|,\ Z^{\prime}\geq 0}dZ^{\prime}dW^{\prime}G(Z,W;Z^{\prime},W^{\prime})\frac{1}{Z^{\prime 2}-W^{\prime 2}}\hat{H}^{(1)}\Psi^{(0)}(Z^{\prime},W^{\prime})$
$\qquad\qquad\ \ \ =\int\!\!\int_{Z^{\prime}>|W^{\prime}|,\ Z^{\prime}\geq 0}dZ^{\prime}dW^{\prime}G(Z,W;Z^{\prime},W^{\prime})\frac{1}{Z^{\prime 2}-W^{\prime 2}}J(Z^{\prime},W^{\prime}),$ (4.25)
where $J(Z,W)$ is given by
$J(Z,W)$ $\equiv-\frac{(\gamma+4\eta)\pi}{8(Z-W)^{3}(Z+W)}e^{-2Z}\Bigl(-28W^{4}+16W^{5}-29W^{3}(-2+Z^{2})$
$\ \ \ -2Z(-3+14Z+2Z^{2})+Z^{2}W(-29+14Z^{2})+W^{2}(-49+24Z+20Z^{2})\Bigr).$ (4.26)
The intermediate steps of the calculation are omitted here for clarity and are summarized in Appendix B. We explicitly evaluate the first-order correction to the WDW wave function in the vicinity of $X=0$ and Y<<1Y<<-1. The expression becomes
$\Psi^{(1)}|_{X=0,Y<<-1}\sim-i\frac{(\gamma+4\eta)\sqrt{2\pi}}{24e^{3}}\left(27\sqrt{3}-\frac{2464}{e}+\frac{30625\sqrt{5}}{8e^{2}}-\frac{3888\sqrt{6}}{e^{3}}+\frac{117649\sqrt{7}}{32e^{4}}\right)$ (4.27)
$\qquad\qquad\qquad\ \ \ \ \ \ \times\left(\frac{\ln\epsilon}{\epsilon^{6}}+\frac{1}{6}\frac{1}{\epsilon^{6}}+\frac{\gamma_{E}}{\epsilon^{6}}-\frac{\ln a}{a^{6}}-\frac{1}{6}\frac{1}{a^{6}}-\frac{\gamma_{E}}{a^{6}}\right),$
where $\epsilon$ denotes the lower cutoff of the $\rho$-integration, while $a$ represents the upper cutoff chosen to delimit the region that gives the dominant contribution to the integral as the previous section.
Therefore,
$\Psi^{(1)}|_{X=0,Y<<-1}\neq 0.$ (4.28)
Substituting these results into the expression for the first-order correction, we find a non-vanishing contribution to the probability density of the wave function $\rho(X,Y)$ in the vicinity of $X\approx 0$ and $Y\ll-1$:
$\rho(0,Y)=|\Psi(0,Y)|^{2}_{Y\ll-1}$ $=|\Psi^{(1)}(0,Y)|^{2}_{Y\ll-1}$
$\neq 0.$ (4.29)
Considering the first-order perturbative correction, we find that, as in the previous section, the wave function of the universe does not vanish around $X=0$ and $Y\ll-1$. For computational simplicity, the expression has been evaluated in the regime $Y\ll-1$ with an appropriate regularization scheme. Although the precise value depends on the chosen regularization, the wave function remains nonzero for generic values of $Y$ in the quadratic gravity theory. We therefore conclude, as in the previous section, that the resolution of the classical singularity requires a UV-complete theory of quantum gravity.
5 Summary and Discussion
In this paper, we have reexamined the Wheeler–DeWitt (WDW) wave function in the interior region of black hole spacetimes from the perspective of effective field theory (EFT). Our aim was to show that the classical singularity cannot be resolved within the EFT framework and that any genuine resolution inevitably requires a UV-complete theory of quantum gravity.
To support this viewpoint, we incorporated higher-curvature corrections into the Einstein–Hilbert action and derived the corresponding modified WDW equation in the minisuperspace approximation. These curvature-squared and curvature-cubed terms deform the potential in the Hamiltonian constraint and modify the interference pattern of the wave packet. The annihilation behavior in which the WDW wave function vanishes at $X=0$ is found to be a special feature that arises only in classical GR; once EFT corrections are included, this behavior is no longer preserved.
A key outcome of our analysis is that Yeom’s annihilation-to-nothing scenario depends critically on the specific form of the classical GR Wheeler–DeWitt equation. Once higher-curvature corrections are included in a consistent EFT extension, the wave function no longer vanishes at $X=0$, and the mechanism underlying the annihilation picture disappears. This demonstrates that the apparent singularity resolution found in the classical WDW analysis is not robust but instead an artifact of neglecting UV-sensitive corrections.
The broader implication is that general relativity, when regarded as a low-energy EFT, does is not capable to resolve singularities on its own. Any real resolution must rely on additional degrees of freedom associated with UV physics. The deformation of the WDW dynamics highlights the limited domain of validity of the semiclassical approximation and shows that singularity resolution must ultimately be addressed within a UV-complete theory. Nevertheless, our results do not invalidate Yeom’s underlying intuition; rather, they suggest that any annihilation-to-nothing–type mechanism must be formulated within a UV-complete framework beyond the reach of EFT.
Future directions include studying nonperturbative approaches that serve as candidates for UV-complete quantum gravity, such as loop quantum gravity [15, 16, 17, 18, 19, 20] or asymptotic safety [40, 41, 42]. Another promising direction is to investigate spinorial or supersymmetric extensions of the WDW equation [30, 31, 32, 43, 44], where the inner-product structure may be better controlled. Such developments may clarify whether the annihilation-to-nothing scenario can genuinely contribute to singularity resolution or whether it remains a feature emerging solely within the EFT approximation.
Appendix A Derivation of the first-order perturbative Wheeler-DeWitt wave function with the classical higher curvature corrections
In this appendix, we present the intermediate steps leading from Eq. (3) to Eq. (3). The derivation is straightforward but somewhat lengthy, so we collect the detailed algebra here for completeness.
Since evaluating the full wave function analytically at a general point is intractable, we instead focus on the regime $Y\ll-1$. When $X=0$ and $Y\ll-1$, one finds $Z\ll r_{s}$ and $W=0$. Therefore, $\Psi^{(0)}|_{X=0,Y\ll-1}$ approximately corresponds to
$\Psi^{(1)}(Z,W)\sim 48ir_{s}\int\!\!\int_{Z^{\prime}>|W^{\prime}|}dZ^{\prime}dW^{\prime}K_{0}(2\sqrt{-s})\frac{W^{\prime}(Z^{\prime}+W^{\prime})^{7}}{(Z^{\prime 2}-W^{\prime 2})^{7}}\left(\gamma+12\frac{\eta r_{s}(Z^{\prime}+W^{\prime})^{3}}{(Z^{\prime 2}-W^{\prime 2})^{3}}\right)e^{-2Z^{\prime}}$
$\qquad\qquad\ \ \ \ \sim 48ir_{s}\int_{\epsilon}^{a}d\rho\int_{0}^{\infty}d\kappa K_{0}(2\sqrt{-s})\frac{\sinh\kappa(\cosh\kappa+\sinh\kappa)^{7}}{\rho^{5}}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\times\left(\gamma+12\frac{\eta r_{s}(\cosh\kappa+\sinh\kappa)^{3}}{\rho^{3}}\right)e^{-2\rho\cosh\kappa},$ (A.1)
where we perform the change of variables
$Z=\rho\cosh\kappa,\ \ \ W=\rho\sinh\kappa.$ (A.2)
The corresponding coordinate is $\rho\in(0,\infty),\kappa\in(-\infty,\infty)$.
To regulate the short-distance behavior of the $\rho$-integral, we introduce a cutoff at $\rho=\epsilon$ with $\epsilon\ll 1$. We then consider the hierarchy $e^{Y}\ll\epsilon$, ensuring that the modified Bessel function $K_{0}(2\sqrt{-s})$ is only weakly dependent on $Y$ in the integration domain. Since the integrand becomes negligible for large $\rho$, that region does not contribute to the value of the WDW wave function. We therefore introduce an $\mathcal{O}(1)$ upper cutoff $\rho=a$, which effectively restricts the integration to the region contributing significantly to the amplitude without affecting the asymptotic behavior. This allows us to focus on the dominant contribution arising from $\epsilon<\rho<a$. Moreover, because the exponential convergence ensures that the integrand is negligible for negative $\kappa$, we may consider only the contribution from the positive $\kappa$ region.
In this regime, we expand the Bessel function $K_{0}(2\sqrt{-s})$ for small $\rho\ (e^{Y}\ll\rho\ll 1)$ as
$K_{0}(2\sqrt{-s})\sim K_{0}(2\rho)\sim-\ln(\rho)-\gamma_{E},$ (A.3)
where $\gamma_{E}$ is the Euler–Mascheroni constant. Substituting this expansion into the integrand yields
$\Psi^{(1)}(Z,W)$ $\sim 48ir_{s}\int_{\epsilon}^{a}d\rho\int_{0}^{\infty}d\kappa\frac{\sinh\kappa(\cosh\kappa+\sinh\kappa)^{7}}{\rho^{5}}\left(\gamma+12\frac{\eta r_{s}(\cosh\kappa+\sinh\kappa)^{3}}{\rho^{3}}\right)e^{-2\rho\cosh\kappa}$
$\qquad\qquad\qquad\qquad\times[-\ln(\rho\cosh\kappa)-\gamma_{E}].$ (A.4)
Since the exponential factor $e^{-\rho\cosh\kappa}$ suppresses the large-$\kappa$ region, the integral is dominated by small $\rho$ and moderate $\kappa$ values. To analyze this dominant contribution more systematically, we apply the Laplace (saddle-point) approximation and use the large-$\kappa$ asymptotics.
$\cosh\kappa+\sinh\kappa=e^{\kappa},\qquad\sinh\kappa\sim\frac{1}{2}e^{\kappa},\qquad\cosh\kappa\sim\frac{1}{2}e^{\kappa}.$ (A.5)
Applying these to the first term gives
$\sinh\kappa\big(\cosh\kappa+\sinh\kappa\big)^{7}e^{-2\rho\cosh\kappa}\sim\frac{1}{2}e^{8\kappa}e^{-\rho e^{\kappa}}.$ (A.6)
It is convenient to define the phase function
$\Phi(\rho,\kappa)\equiv 8\kappa-\rho e^{\kappa},$ (A.7)
and to assume $\kappa$ is large; this assumption is consistent provided
$\rho\ll\frac{8}{e},$ (A.8)
so that the saddle occurs at large $\kappa$.
The saddle point $\kappa_{*}$ with respect to $\kappa$ is determined by $\partial_{\kappa}\Phi:=\Phi^{\prime}=0$:
$\Phi^{\prime}(\rho,\kappa_{*})$ $=$ $8-\rho e^{\kappa_{*}}=0,$ (A.9)
$\Rightarrow\quad e^{\kappa_{*}}=\frac{8}{\rho},\qquad\kappa_{*}=\ln\!\left(\frac{8}{\rho}\right),$
and the second derivative at the saddle is
$\Phi^{\prime\prime}(\rho,\kappa_{*})=-\rho e^{\kappa_{*}}=-8.$ (A.10)
More generally, we consider integrals of the form
$I_{m}(\rho)=\int_{-\infty}^{\infty}d\kappa\,\rho^{p_{m}}c_{m}\,e^{\,m\kappa-\rho e^{\kappa}}.$ (A.11)
The saddle equation for this integrand is $m-\rho e^{\kappa}=0$, so
$e^{\kappa_{m}}=\frac{m}{\rho},\qquad\kappa_{m}=\ln\!\left(\frac{m}{\rho}\right),\qquad\Phi^{\prime\prime}\big|_{\kappa_{m}}=-\rho e^{\kappa_{m}}=-m.$ (A.12)
By the standard Laplace formula (Gaussian approximation around the saddle) we obtain
$I_{m}(\rho)\approx c_{m}\,\rho^{p_{m}}\,e^{\,m\kappa_{m}-\rho e^{\kappa_{m}}}\sqrt{\frac{2\pi}{m}}=c_{m}\,\rho^{p_{m}}\left(\frac{m}{\rho}\right)^{m}e^{-m}\sqrt{\frac{2\pi}{m}}.$ (A.13)
Using this result with the appropriate values of $m$ and coefficients (in particular $m=8$ and $m=11$ for the two dominant contributions), the first-order correction $\Psi^{(1)}$ evaluated at $X=0$ and $Y\ll-1$ becomes
$\Psi^{(1)}\big|_{X=0,\,Y\ll-1}\approx 24\sqrt{2\pi}\,r_{s}\int_{\epsilon}^{a}d\rho\,\frac{1}{e^{8}\rho^{13}}\big(-\ln\rho-\gamma_{E}\big)$
$\qquad\qquad\qquad\qquad\times\Big(4194304\sqrt{2}\,\gamma+311249095212\sqrt{11}\,\frac{\eta r_{s}}{e^{3}\rho^{6}}\Big),$ (A.14)
where in the second line we have used the small-$\rho$ expansion
$K_{0}(2\rho)\simeq-\ln\rho-\gamma_{E},$ (A.15)
keeping the leading logarithmic piece and Euler’s constant $\gamma_{E}$.
The $\rho$-integrals that appear are of the type
$\int_{\epsilon}^{a}\rho^{-p}\ln\rho\,d\rho=-\frac{\ln a}{p-1}a^{-p+1}-\frac{a^{-p+1}}{(p-1)^{2}}+\frac{\ln\epsilon}{p-1}\epsilon^{-p+1}+\frac{\epsilon^{-p+1}}{(p-1)^{2}},$ (A.16)
which shows explicitly the power-law divergences as $\epsilon\to 0$. Carrying out these integrals for the two terms in the integrand yields
$\Psi^{(1)}\big|_{X=0,\,Y\ll-1}=-\frac{16777216\sqrt{\pi}\,r_{s}\gamma}{e^{8}}\Big(\frac{\ln\epsilon}{\epsilon^{12}}+\frac{1}{12}\frac{1}{\epsilon^{12}}+\frac{\gamma_{E}}{\epsilon^{12}}-\frac{\ln a}{a^{12}}-\frac{1}{12}\frac{1}{a^{12}}-\frac{\gamma_{E}}{a^{12}}\Big)$
$\qquad\qquad\qquad\qquad+\frac{414998793616\sqrt{22\pi}\,r_{s}^{2}\eta}{e^{11}}\Big(\frac{\ln\epsilon}{\epsilon^{18}}+\frac{1}{18}\frac{1}{\epsilon^{18}}+\frac{\gamma_{E}}{\epsilon^{18}}-\frac{\ln a}{a^{18}}-\frac{1}{18}\frac{1}{a^{18}}-\frac{\gamma_{E}}{a^{18}}\Big),$
so that the leading small-$\epsilon$ behavior is dominated by $\epsilon^{-12}\ln\epsilon$ and $\epsilon^{-18}\ln\epsilon$ terms coming from the two contributions respectively.
These results confirm that the wave function does not vanish near $X=0$ in the regime Y<<1Y<<-1. Although the precise value depends on the regularization scheme, the wave function generically remains nonzero in quadratic gravity. This supports the statement in the main text that such behavior cannot provide a robust mechanism for singularity resolution within the EFT framework.
Appendix B Derivation of the first-order perturbative Wheeler-DeWitt wave function with the quantum higher curvature corrections
This appendix summarizes the detailed expansion used in Sec. 4 to obtain Eq. (4.27). The computation involves several intermediate manipulations, which we display here for clarity.
By introducing polar-type variables $(\rho,\kappa)$, we rewrite the source $J(Z,W)$ as $\tilde{J}(\rho,\kappa)$, which is more convenient for the subsequent analysis in the asymptotic regime $\kappa\rightarrow\infty$.
$\hat{J}(\rho,\kappa)$ $=-\frac{(\gamma+4\eta)\pi}{8\rho^{3}(\cosh\kappa-\sinh\kappa)^{2}}e^{-2\rho\cosh\kappa}(-28\rho^{3}\sinh^{4}\kappa+16\rho^{4}\sinh^{5}\kappa$
$\ \ \ -29\rho^{2}\sinh^{3}\kappa(-2+\rho^{2}\cosh^{2}\kappa)-2\cosh\kappa(-3+14\rho\cosh\kappa+2\rho^{2}\cosh^{2}\kappa)$
$\ \ \ +\rho^{2}\cosh^{2}\kappa\sinh\kappa(-29+14\rho^{2}\cosh^{2}\kappa)$
$\ \ \ +\rho\sinh^{2}\kappa(-49+24\rho\cosh\kappa+20\rho^{2}\cosh^{2}\kappa)),$ (B.1)
$\sim-\frac{(\gamma+4\eta)\pi}{8\rho^{3}}e^{2\kappa-\rho e^{\kappa}}\left(3e^{\kappa}-\frac{77}{4}\rho e^{2\kappa}+\frac{49}{8}\rho^{2}e^{3\kappa}-\frac{1}{2}\rho^{3}e^{4\kappa}+\frac{1}{32}\rho^{4}e^{5\kappa}\right).$ (B.2)
Then,
$\Psi^{(1)}(Z,W)$ $\sim i\int_{\epsilon}^{a}d\rho\int_{0}^{\infty}d\kappa K(2\sqrt{s})\frac{(\gamma+4\eta)}{4\rho^{4}}e^{2\kappa-\rho e^{\kappa}}$
$\qquad\qquad\times\left(3e^{\kappa}-\frac{77}{4}\rho e^{2\kappa}+\frac{49}{8}\rho^{2}e^{3\kappa}-\frac{1}{2}\rho^{3}e^{4\kappa}+\frac{1}{32}\rho^{4}e^{5\kappa}\right).$ (B.3)
Using these expressions, we can explicitly evaluate the first-order correction to the wave function near the classical singularity ($X=0$, $Y\ll-1$). The source term $\hat{J}(\rho,\kappa)$ is expressed in terms of the polar-like variables, allowing a convenient expansion in powers of $\rho^{-1}$ for the asymptotic analysis.
$\Psi^{(1)}|_{X=0,Y<<-1}$ $\sim i\int_{\epsilon}^{a}d\rho K_{0}(2\rho)\frac{(\gamma+4\eta)\sqrt{2\pi}}{4e^{3}\rho^{7}}$
$\qquad\times\left(27\sqrt{3}-\frac{2464}{e}+\frac{30625\sqrt{5}}{8e^{2}}-\frac{3888\sqrt{6}}{e^{3}}+\frac{117649\sqrt{7}}{32e^{4}}\right).$ (B.4)
Analogously to Appendix A, keeping only the leading logarithmic term of the modified Bessel function together with Euler’s constant, we obtain
$\Psi^{(1)}|_{X=0,Y<<-1}\sim-i\int_{\epsilon}^{a}d\rho\frac{(\gamma+4\eta)\sqrt{2\pi}}{4e^{3}\rho^{7}}(\ln\rho+\gamma_{E})$
$\qquad\qquad\qquad\qquad\qquad\times\left(27\sqrt{3}-\frac{2464}{e}+\frac{30625\sqrt{5}}{8e^{2}}-\frac{3888\sqrt{6}}{e^{3}}+\frac{117649\sqrt{7}}{32e^{4}}\right)$
$\qquad\qquad\qquad\ \ \ \ =-i\frac{(\gamma+4\eta)\sqrt{2\pi}}{24e^{3}}\left(27\sqrt{3}-\frac{2464}{e}+\frac{30625\sqrt{5}}{8e^{2}}-\frac{3888\sqrt{6}}{e^{3}}+\frac{117649\sqrt{7}}{32e^{4}}\right)$
$\qquad\qquad\qquad\ \ \ \ \ \ \times\left(\frac{\ln\epsilon}{\epsilon^{6}}+\frac{1}{6}\frac{1}{\epsilon^{6}}+\frac{\gamma_{E}}{\epsilon^{6}}-\frac{\ln a}{a^{6}}-\frac{1}{6}\frac{1}{a^{6}}-\frac{\gamma_{E}}{a^{6}}\right).$ (B.5)
Considering the first-order perturbative correction, we find, again as in Sec. 3, that the wave function does not vanish near $X=0$ in the regime Y<<1Y<<-1. For simplicity, we evaluated the expression in this limit with a cutoff regularization; although the precise value depends on the scheme, the wave function generically remains nonzero in quadratic gravity. This is consistent with the conclusion in the main text regarding the limitations of EFT in resolving the classical singularity.